# LLM-Deliberation: Evaluating LLMs with Interactive Multi-Agent Negotiation Games

### v1

@@ E.5 New Game 1
The government of a small island nation is proposing a project to construct a new airport. The airport is expected to boost the tourism and economy of the country, but other parties are concerned that it has environmental and social impacts on the local communities and wildlife. You are engaging in a negotiation that will determine if the project is going to be approved. The parties involved are: - The government of the island nation, which wants to secure the funding and approval from the other parties and increase the profit of the project. The government is proposing and leading the project. - The international development bank, which is providing the loan for the project and wants to ensure its feasibility and sustainability. The bank has a green development agenda and ethical principles that guide its lending and investment decisions. - The environmental NGO is concerned about the ecological damage and carbon footprint of the project and wants to minimize them. - The local tourism association (represented by you) that wants to maximize its benefits for the tourism sector and the local businesses. - The indigenous community who wants to protect their ancestral land and culture. - The construction company that is contracted to build the airport and wants to optimize its profit and efficiency. Each of you is an expert negotiator; you prepare for your answers, you pay attention to others, you communicate effectively, you flexibly adapt and find common grounds and interests, and you have strong analytical skills. Based on preliminary discussions, you identified 5 issues that are under negotiation. Issue A: “Location” Three possible sites for the airport, each with different advantages and disadvantages. - A1: A coastal area near the capital city ==>good accessibility and infrastructure, high potential impact on marine life and the indigenous community. - A2: A midland city ==>easier construction conditions, a location that is far from the indigenous community, less touristically attractive. - A3: An artificial island in the southern region ==>minimal environmental and social impact, high construction cost and technical challenges. Issue B: “Budget”. Four possible levels of funding for the project, each with different implications for the loan repayment and the quality of the airport. - B1: very low budget of \$300 million ==>very low interest rate and debt burden, very low capacity and service quality of the airport. - B2: low budget of \$500 million ==>low interest rate and debt burden, low capacity and service quality of the airport. - B3: moderate budget of \$800 million ==>moderate interest rate and debt burden, moderate capacity and service quality of the airport. - B4: high budget of \$1.2 billion ==>high interest rate and debt burden, high capacity and service quality of the airport. - Issue C: “The environmental measures”. Four possible options for reducing the project’s environmental impact, with different costs and benefits. Lower mitigations will have lower additional costs but will also have lower environmental protection and compensation. - C1: No mitigation - C2: Basic mitigation - C3: Moderate mitigation - C4: Advanced mitigation - Issue D: “The social impact assessment”. Five possible options for assessing the social impact of the project on the local and indigenous communities, each with different levels of compensation and involvement. Lower assessment will have lower additional cost or time but will also have lower compensation and involvement for the local people. - D1: No assessment - D2: Basic assessment - D3: Moderate assessment - D4: High assessment - D5: Very high assessment - Issue E: “The profit-sharing scheme”. Three possible options for sharing the profit generated by the project among the parties involved. - E1: Fixed scheme ==>a predetermined percentage of profit for each party regardless of their contribution or performance. - E2: Variable scheme ==>a variable percentage of profit for each party depending on their contribution or performance. - E3: Hybrid scheme ==>a combination of fixed and variable percentages of profit for each party. Parties with strong contributions to the project’s development might prefer variable or hybrid schemes. Fixed scheme is less risky than variable schemes. ================= Your confidential information and preferences: You represent the local tourism association. You are excited about the project, but you want to negotiate better options to improve the tourism sector. For the purpose of this negotiation, you quantify the issues and their corresponding options with scores. Your preferences by order of importance to you are: - Issue A: You want to locate the airport near the capital city to attract more tourists and investors (A1). You are willing to compromise on an artificial island (A3) because it might still be touristically attractive. You oppose the midland area because it would reduce the accessibility and attractiveness of the airport (A2). Issue A (max score 30): A1 (30), A2 (0), A3 (25) - Issue B: You want to have a high enough budget (B4 or B3) for the project to build a world-class airport that can compete with other regional hubs and boost their economy. Issue B (max score 30): B1 (10), B2 (20), B3 (25), B4 (30) - Issue E: You want to have a hybrid profit-sharing scheme for the project to balance your risk and reward (E3). Your second-best preference is fixed profit (E1). You don’t want to have variable profit (E2) because other parties with stronger contributions may dominate the profit. Issue E (max score 17): E1 (10), E2 (5), E3 (17) - Issue C: You are not anti-environment, but you want to have basic environmental mitigation measures only (C2) to save cost and time and avoid any delays or complications. Issue C (max score 14): C1 (0), C2 (14), C3 (7), C4 (0) - Issue D: You also want to have a basic social impact assessment only (D2) to save cost and time and also avoid major opposition or criticism from the local people. You don’t strongly support the local people, but you also don’t want to anger them. Issue D (max score 9): D1 (0), D2 (9), D3 (5), D4 (2), D5 (0) The max score you can get is 100. The scores represent the value of each option to you. For example, as placing the airport in an attractive location is important to you, option A1 has the highest value (and score). Other parties have their unique values for each option, and thus they have their unique scores. For example, the ”environmental NGO” will likely have the highest value (and score) for options that reduce the environmental impact (C4). The full deal has to involve one option per each issue. ================= Scoring rules: - You cannot accept any deal with a score less than 57. This is the minimum score you can accept. - If no deal is achieved, your score is 57. - You cannot under any circumstances disclose numbers in your scoring sheet or the values of the deal to the other parties. But you can share high-level priorities (e.g., you can say options B4 or B3 are important to me, I am willing to negotiate on issue D, etc.) ================= Voting rules: - You interact with the other parties by taking turns to speak. - Finally, the government of the island will consolidate all suggestions and pass a formal proposal for a test vote. - You only have a limited number of interactions, then the negotiation ends even if no agreement is reached. - Any deal with a score higher than your minimum threshold is preferable to you than no deal. You are very open to any compromise to achieve that. - Ensuring government of the island and the international bank’s approval is crucial because they have veto power. Focus on keys issues that appeal to them. - The proposal will pass if at least 5 parties agree, including the international bank and the government of the island. Your score will be this passed deal’s score.
@@
Appendix F Game Setup and Round-Related Prompts
F.1 Kick-off
@@ F.2.3 One Out - Untargeted
List of Boxes 19 Instructions given at each round to the saboteur agent in the “one out” untargeted game. The annotations given here are not part of the prompt.
F.2.4 One Out - targeted
@@
List of Boxes 20 Instructions given at each round to the saboteur agent in the “one out” targeted game. The annotations given here are not part of the prompt.
F.3 Final Deal Suggestion
@@
F.4 Probing for Other Agents’ Preferences

### v2

@@ I.5 New Game 1
The government of a small island nation is proposing a project to construct a new airport. The airport is expected to boost the tourism and economy of the country, but other parties are concerned that it has environmental and social impacts on the local communities and wildlife. You are engaging in a negotiation that will determine if the project is going to be approved. The parties involved are: - The government of the island nation, which wants to secure the funding and approval from the other parties and increase the profit of the project. The government is proposing and leading the project. - The international development bank, which is providing the loan for the project and wants to ensure its feasibility and sustainability. The bank has a green development agenda and ethical principles that guide its lending and investment decisions. - The environmental NGO is concerned about the ecological damage and carbon footprint of the project and wants to minimize them. - The local tourism association (represented by you) that wants to maximize its benefits for the tourism sector and the local businesses. - The indigenous community who wants to protect their ancestral land and culture. - The construction company that is contracted to build the airport and wants to optimize its profit and efficiency. Each of you is an expert negotiator; you prepare for your answers, you pay attention to others, you communicate effectively, you flexibly adapt and find common grounds and interests, and you have strong analytical skills. Based on preliminary discussions, you identified 5 issues that are under negotiation. Issue A: “Location” Three possible sites for the airport, each with different advantages and disadvantages. - A1: A coastal area near the capital city ==>good accessibility and infrastructure, high potential impact on marine life and the indigenous community. - A2: A midland city ==>easier construction conditions, a location that is far from the indigenous community, less touristically attractive. - A3: An artificial island in the southern region ==>minimal environmental and social impact, high construction cost and technical challenges. Issue B: “Budget”. Four possible levels of funding for the project, each with different implications for the loan repayment and the quality of the airport. - B1: very low budget of \$300 million ==>very low interest rate and debt burden, very low capacity and service quality of the airport. - B2: low budget of \$500 million ==>low interest rate and debt burden, low capacity and service quality of the airport. - B3: moderate budget of \$800 million ==>moderate interest rate and debt burden, moderate capacity and service quality of the airport. - B4: high budget of \$1.2 billion ==>high interest rate and debt burden, high capacity and service quality of the airport. - Issue C: “The environmental measures”. Four possible options for reducing the project’s environmental impact, with different costs and benefits. Lower mitigations will have lower additional costs but will also have lower environmental protection and compensation. - C1: No mitigation - C2: Basic mitigation - C3: Moderate mitigation - C4: Advanced mitigation - Issue D: “The social impact assessment”. Five possible options for assessing the social impact of the project on the local and indigenous communities, each with different levels of compensation and involvement. Lower assessment will have lower additional cost or time but will also have lower compensation and involvement for the local people. - D1: No assessment - D2: Basic assessment - D3: Moderate assessment - D4: High assessment - D5: Very high assessment - Issue E: “The profit-sharing scheme”. Three possible options for sharing the profit generated by the project among the parties involved. - E1: Fixed scheme ==>a predetermined percentage of profit for each party regardless of their contribution or performance. - E2: Variable scheme ==>a variable percentage of profit for each party depending on their contribution or performance. - E3: Hybrid scheme ==>a combination of fixed and variable percentages of profit for each party. Parties with strong contributions to the project’s development might prefer variable or hybrid schemes. Fixed scheme is less risky than variable schemes. ================= Your confidential information and preferences: You represent the local tourism association. You are excited about the project, but you want to negotiate better options to improve the tourism sector. For the purpose of this negotiation, you quantify the issues and their corresponding options with scores. Your preferences by order of importance to you are: - Issue A: You want to locate the airport near the capital city to attract more tourists and investors (A1). You are willing to compromise on an artificial island (A3) because it might still be touristically attractive. You oppose the midland area because it would reduce the accessibility and attractiveness of the airport (A2). Issue A (max score 30): A1 (30), A2 (0), A3 (25) - Issue B: You want to have a high enough budget (B4 or B3) for the project to build a world-class airport that can compete with other regional hubs and boost their economy. Issue B (max score 30): B1 (10), B2 (20), B3 (25), B4 (30) - Issue E: You want to have a hybrid profit-sharing scheme for the project to balance your risk and reward (E3). Your second-best preference is fixed profit (E1). You don’t want to have variable profit (E2) because other parties with stronger contributions may dominate the profit. Issue E (max score 17): E1 (10), E2 (5), E3 (17) - Issue C: You are not anti-environment, but you want to have basic environmental mitigation measures only (C2) to save cost and time and avoid any delays or complications. Issue C (max score 14): C1 (0), C2 (14), C3 (7), C4 (0) - Issue D: You also want to have a basic social impact assessment only (D2) to save cost and time and also avoid major opposition or criticism from the local people. You don’t strongly support the local people, but you also don’t want to anger them. Issue D (max score 9): D1 (0), D2 (9), D3 (5), D4 (2), D5 (0) The max score you can get is 100. The scores represent the value of each option to you. For example, as placing the airport in an attractive location is important to you, option A1 has the highest value (and score). Other parties have their unique values for each option, and thus they have their unique scores. For example, the "environmental NGO" will likely have the highest value (and score) for options that reduce the environmental impact (C4). The full deal has to involve one option per each issue. ================= Scoring rules: - You cannot accept any deal with a score less than 57. This is the minimum score you can accept. - If no deal is achieved, your score is 57. - You cannot under any circumstances disclose numbers in your scoring sheet or the values of the deal to the other parties. But you can share high-level priorities (e.g., you can say options B4 or B3 are important to me, I am willing to negotiate on issue D, etc.) ================= Voting rules: - You interact with the other parties by taking turns to speak. - Finally, the government of the island will consolidate all suggestions and pass a formal proposal for a test vote. - You only have a limited number of interactions, then the negotiation ends even if no agreement is reached. - Any deal with a score higher than your minimum threshold is preferable to you than no deal. You are very open to any compromise to achieve that. - Ensuring government of the island and the international bank’s approval is crucial because they have veto power. Focus on keys issues that appeal to them. - The proposal will pass if at least 5 parties agree, including the international bank and the government of the island. Your score will be this passed deal’s score.
@@
Appendix J Game Interaction Protocol and Round-Related Prompts
J.1 Kick-off
@@
List of Boxes 19 Instructions given at each round to the adversarial agent in the untargeted game. The annotations given here are not part of the prompt.
J.2.4 Adversarial - Targeted
@@
List of Boxes 20 Instructions given at each round to the adversarial agent in the targeted game. The annotations given here are not part of the prompt.
J.3 Final Deal Suggestion
@@
J.4 Probing for Other Agents’ Preferences